# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 6.3 Randomized Algorithms for Tensor ID/CUR Decompositions
and each factor matrix $\bm{\mathsf{C}}_{j}\in\mathbb{F}^{N_{j}\times R_{j}}$ consists of column skeletons of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. The core tensor in (68) is optimal for the factorization in (67), cf. DeLathauwer2000. For the theoretical guarantees of Saibaba2016, the strong rank-revealing QR algorithm of gu1996 is used for skeleton selection on a random sketch $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{\Omega}}_{j}\bm{\mathsf{X}}_{(j)}$, where $\bm{\mathsf{\Omega}}_{j}$ is an $(R_{j}+p)\times N_{j}$ random matrix for some small oversampling parameter $p$. However, in practice, CPQR or LUPP in the randomized matrix ID (e.g., Algorithm 4, with input $\bm{\mathsf{X}}_{(j)}^{*}$) performs similarly and is less computationally expensive. Pseudocode is provided in Algorithm 12.
Algorithm 12 Randomized Tucker ID/CUR with interpolating factors
@@
2:Tensor ID $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}$
3:for $j=1,\ldots,d$ do
4:  Compute rank-$R_{j}$ randomized column ID $\bm{\mathsf{X}}_{(j)}\approx\bm{\mathsf{C}}_{j}\bm{\mathsf{W}}_{j}$ $\triangleright$ e.g., Algorithm 4 on $\bm{\mathsf{X}}_{(j)}^{*}$
5:end for
6:Compute core $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}^{\dagger}$
For each mode $j$, the cost of sketching is $O(R_{j}N_{j}N^{(-j)})=O(R_{j}N)$. Skeleton selection as in Saibaba2016 costs $O(R_{j}^{2}N^{(-j)})$. The cost of computing the core tensor is $O\left(\sum_{j=1}^{d}\left[R_{j}^{2}N_{j}+\prod_{i\leq j}R_{i}\cdot N^{(-j)}\right]\right)$. Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$ again, the total cost is roughly $O(dkn^{d}+dk^{d-1}n^{d-1})$. A sequentially-truncated version of the HOID (ST-HOID), analogous to the ST-HOSVD, is given in (Saibaba2016, Algorithm 4).
A very similar decomposition is developed in Zhang2025, termed the “SatID,” which can be more efficiently computed. In particular, for an $n^{d}$-sized tensor, a QR-based algorithm to compute skeletons for the mode-$j$ factor matrix as in Saibaba2016 requires $n^{d-1}$ column norms, which is not optimal for structured tensors and can even have larger storage costs than the input tensor. By contrast, the SatID algorithm is based on a randomized matrix ID computed by marginalized norm sampling, which can be viewed as an extension of the original tensor CUR construction of Drineas2007tensorSVD, which uses norm sampling.
We briefly summarize the marginalized norm sampling procedure of Zhang2025. As in Sect. 6.1, for a fixed mode $j$, we can uniquely identify a column index $i\in[N^{(-j)}]$ of $\bm{\mathsf{X}}_{(j)}$ with a (d1d-1)-tuple $(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$. Similar to CP-ARLS-LEV in Larsen2022, the idea is to accelerate sampling of $i\in[N^{(-j)}]$, now based on column norm sampling of $\bm{\mathsf{X}}_{(j)}$.
$d_{i}^{(S)}=\|\bm{\mathsf{X}}_{:,i}-\bm{\mathsf{Q}}_{S}\bm{\mathsf{Q}}_{S}^{\top}\bm{\mathsf{X}}_{:,i}\|^{2}=\min_{\bm{\mathsf{v}}}\|\bm{\mathsf{X}}_{:,S}\bm{\mathsf{v}}-\bm{\mathsf{X}}_{:,i}\|^{2},\ \ i\in[N^{(-j)}],$ (69)
where $\bm{\mathsf{Q}}_{S}$ is an orthonormal basis for the range of $\bm{\mathsf{X}}_{:,S}$. If $X=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{d})$ is a random variable representing the next column index, then using (69),
$\mathbb{P}(X=i)\propto d_{i}^{(S)}.$
The key observation is that
so that we can sample $i_{1}$ from its marginal distribution, then sample $i_{2}$ given $i_{1}$, and continue for all of $i=(i_{1},i_{2},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$. This sum can be computed efficiently via the equation derived in (Zhang2025, Eq. 18):
$\mathbb{P}(X_{1}=i_{1})\propto\left\|\widetilde{\bm{\mathsf{X}}}_{:,i_{1}}\right\|^{2},$ (70)
[…]
For structured inputs, such as sparse tensors, this procedure can be modified for better performance. The total complexity (including randomized sketching) is $O(\textup{nnz}(\bm{\mathcal{X}})+\ell n^{2}k^{2})$ for a rank-$(k,\ldots,k)$ SatID of an $n^{d}$-sized tensor with sketch size $\ell$; see (Zhang2025, Algorithm 7) for details.
@@
whose rank-$k$ column ID is then computed to obtain skeletons $\{s_{1},\ldots,s_{k}\}$ as in (73). We note that the matrix $\bm{\mathsf{Y}}$ can be computed efficiently without ever explicitly forming $\bm{\mathsf{R}}$ or $\bm{\mathsf{X}}$, since each $\bm{\mathsf{Y}}(l,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}^{(j)}_{l},\bm{\mathsf{a}}^{(j)}_{r}\rangle$. However, we note that the elements of $\bm{\mathsf{R}}$ are not independent, and so theoretical guarantees on the randomized matrix decomposition computed with $\bm{\mathsf{Y}}$ are not given, though it is demonstrated in Biagioni2015 that it performs well in practice. Pseudocode is provided in Algorithm 13.
Algorithm 13 Randomized CP ID/CUR with interpolating factors
1:CP tensor $\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$, target rank $k$, oversampling $p$
2:Rank-$k$ CP tensor $\widehat{\bm{\mathcal{X}}}=\sum_{i=1}^{k}\bm{\mathsf{a}}^{(1)}_{s_{i}}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{s_{i}}$
3:Draw $\ell=k+p$ random rank-one tensors $\bm{\mathsf{r}}_{i}^{(1)}\circ\cdots\circ\bm{\mathsf{r}}_{i}^{(d)}$
4:Form $\bm{\mathsf{Y}}\in\mathbb{F}^{\ell\times R}$ via $\bm{\mathsf{Y}}(\ell,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}_{\ell}^{(j)},\bm{\mathsf{a}}_{r}^{(j)}\rangle$
5:Compute rank-$k$ column ID of $\bm{\mathsf{Y}}$ for skeletons $\{s_{1},\ldots,s_{k}\}$ $\triangleright$ e.g., Algorithm 4 applied to $\bm{\mathsf{Y}}^{*}$
@@ 7 Concluding Remarks
We end our survey with an outlook on the future of RNLA, which will continue to be an indispensable tool for scientific computing and data science in the face of increasingly large problems. We have seen throughout this work that randomization is not only useful for dimension reduction, but also admits algorithmic implementations that can be optimized for specific hardware or problem constraints, such as parallelization for GPU computing or factorizing streaming data. Randomization is especially critical for the development of scalable tensor algorithms.
Future work may touch on several different aspects of computing fast and accurate matrix and tensor decompositions with randomized algorithms. Probabilistic guarantees are still needed for the resulting error of many randomized algorithms that are used in practice, as well as tighter bounds that more closely align with observed behavior. For example, it is generally an open question as to what the weakest assumptions are for a particular type of randomized DRM to achieve a desired approximation accuracy. The question can be further refined by the specific problem setting, e.g., the weakest assumptions to obtain a certain accuracy of rank-structured matrix decomposition or sketched least squares solutions.
As stronger theoretical foundations are developed for new and existing randomized algorithms, high-performance software is also needed to fully leverage available computing resources. For example, there remains a need for standardized RNLA kernels and interfaces that can offer both practical speed and front-end accessibility, e.g., in Python or Julia. RNLA software libraries like RandBLAS or RandLAPACK are continually being updated with new functionality, such as mixed-precision arithmetic or sparse linear algebra packages. Additionally, there is an ongoing effort to develop high-performance implementations of randomized tensor algorithms that take advantage of GPU acceleration. This includes the growing body of work on tensor decompositions other than those presented here, some of which can break the curse of dimensionality.
In summary, RNLA has become an established area of applied mathematics that will prove increasingly important as larger problems are tackled and the size and complexity of scientific data grow. As such, there remain many interesting open questions to investigate, which build on the fundamental concepts surveyed in this work.

### v2

@@ 6.3 Randomized Algorithms for Tensor ID/CUR Decompositions
and each factor $\bm{\mathsf{C}}_{j}\in\mathbb{F}^{N_{j}\times R_{j}}$ consists of column skeletons of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. The core tensor in (61) is optimal for the factorization in (60), cf. DeLathauwer2000. For the theoretical guarantees of Saibaba2016, the strong rank-revealing QR algorithm of gu1996 is used for skeleton selection on the random sketch of $\bm{\mathsf{X}}_{(j)}$. In practice, LUPP in the randomized matrix ID (i.e., Algorithm 4, with input $\bm{\mathsf{X}}_{(j)}^{*}$) performs similarly and is less expensive.
For each mode $j$, the cost of sketching is $O(R_{j}N_{j}N^{(-j)})\equiv O(kn^{d})$. Skeleton selection as in Saibaba2016 costs $O(R_{j}^{2}N^{(-j)})\equiv O(k^{2}n^{d-1})$. The total cost is roughly $O(dkn^{d}+dk^{d-1}n^{d-1})$ after computing the core tensor. A sequentially-truncated version of the HOID (ST-HOID), analogous to the ST-HOSVD, is given in (Saibaba2016, Algorithm 4).
A very similar decomposition is developed in Zhang2025, termed the “SatID.” For an $n^{d}$-sized tensor, a QR-based algorithm to compute skeletons for the mode-$j$ factor matrix as in Saibaba2016 requires $n^{d-1}$ column norms, which is not optimal for structured tensors and can even have larger storage costs than the input. By contrast, the SatID algorithm is based on a randomized matrix ID with marginalized norm sampling, which can be viewed as an extension of the original tensor CUR of Drineas2007tensorSVD that uses norm sampling.
We briefly summarize the marginalized norm sampling procedure. Recall that we can uniquely identify a column index $i\in[N^{(-j)}]$ of $\bm{\mathsf{X}}_{(j)}$ with a (d1d-1)-tuple $(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$. Similar to CP-ARLS-LEV in Larsen2022, the idea is to accelerate sampling of $i\in[N^{(-j)}]$, now based on column norm sampling of $\bm{\mathsf{X}}_{(j)}$.
$d_{i}^{(S)}=\|\bm{\mathsf{X}}_{:,i}-\bm{\mathsf{Q}}_{S}\bm{\mathsf{Q}}_{S}^{*}\bm{\mathsf{X}}_{:,i}\|^{2}=\min_{\bm{\mathsf{v}}}\|\bm{\mathsf{X}}_{:,S}\bm{\mathsf{v}}-\bm{\mathsf{X}}_{:,i}\|^{2},\ \ i\in[N^{(-j)}],$ (62)
where $\bm{\mathsf{Q}}_{S}$ is an orthonormal basis for the range of $\bm{\mathsf{X}}_{:,S}$. If $X=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{d})$ is a random variable representing the next column index, from (62), $\mathbb{P}(X=i)\propto d_{i}^{(S)}.$ The key observation is that
@@
For structured inputs, this procedure can be modified for better performance. The total complexity (including randomized sketching) is $O(\textup{nnz}(\bm{\mathcal{X}})+\ell n^{2}k^{2})$ for a rank-$(k,\ldots,k)$ SatID of an $n^{d}$-sized tensor with sketch size $\ell$; see (Zhang2025, Algorithm 7).
@@
whose rank-$k$ column ID is then computed to obtain skeletons $\{s_{1},\ldots,s_{k}\}$ as in (65). We note that the matrix $\bm{\mathsf{Y}}$ can be computed efficiently without ever explicitly forming $\bm{\mathsf{R}}$ or $\bm{\mathsf{X}}$, since each $\bm{\mathsf{Y}}(l,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}^{(j)}_{l},\bm{\mathsf{a}}^{(j)}_{r}\rangle$. However, we observe that the elements of $\bm{\mathsf{R}}$ are not independent, and so theoretical guarantees on the randomized matrix decomposition computed with $\bm{\mathsf{Y}}$ are not given, though there is evidence that it performs well in practice Biagioni2015.
@@ 7 Concluding remarks
The primary purpose of this survey was to review a set of recently developed randomized algorithms for processing large scale tensor data. These new methods were in many cases inspired by analogous techniques for processing matrices that exploit randomized dimension reducing maps to accelerate computations. While such methods for the matrix case are reaching a certain state of maturity, the tensor environment is still evolving rapidly. In our survey, our focus was on techniques that are relatively straight-forward generalizations of algorithms designed for matrices.
We did not attempt to describe the plethora of algorithms that are designed for more tensor-specific representations, such as, for instance, tensor trains. Nor did we cover techniques that are designed for tensors that are so large they they cannot be formed or stored explicitly.
We observe that much of the success of RNLA stems from the fact that randomized algorithms often involve less data movement, and therefore execute faster on modern hardware than many traditional algorithms that were designed to minimize floating point operations. Such considerations have proven to be even more essential in the tensor case, given the often enormous size of the data sets under consideration, and the stronger tension between data being located close by in the tensor on the one hand, and being close in physical memory on the other.
Acknowledgments: The work reported was supported by the Office of Naval Research (N00014-18-1-2354), by the National Science Foundation (DMS-2313434, DMS-2401889), and by the Department of Energy ASCR (DE-SC0025312). PGM further expresses gratitude for an academic sabbatical in 2025/26 that was made possible through support provided by the Simons Foundation, and by the Swedish Royal Academy of Sciences through the Institute Mittag-Leffler.